OpenStudy (anonymous):

how to get the derivative of e^x^2+e^4x ?

OpenStudy (anonymous):

I'm not sure if I write the function correctly? \[\large (e^{x})^{2} + e^{4x}\]

OpenStudy (anonymous):

\[e^{2x} + e^{4x}\] \[= 2e^{2x} + 4e^{4x}\] \[= 2( 2e^{4x} + e^{2x})\] Is this correct? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

Correct! rewrite the chain rule that you used

OpenStudy (anonymous):

The rule is \[\large \frac{ d }{ dx } e^{f(x)} = \frac{ d }{ dx } f(x) * e^{f(x)}\]

OpenStudy (anonymous):

Thank you very much, xapproach! :))

OpenStudy (xapproachesinfinity):

Anytime!

zepdrix (zepdrix):

There's a difference between: \(\Large\rm (e^x)^2\) and \(\Large\rm e^{x^2}\) I'm not sure which one she meant, hopefully you have the correct one heh c:

OpenStudy (xapproachesinfinity):

yeah! I was not sure what he wrote too was it exp(x)2 or epx(x^2)

OpenStudy (xapproachesinfinity):

only the one who posted knows what is that.

OpenStudy (anonymous):

So can we try \[\large e^{x^{2}}\] :)) a little harder

zepdrix (zepdrix):

Hmm yah that one could be fun c:

zepdrix (zepdrix):

Give it a shot stud! :3

OpenStudy (anonymous):

\[\large = 2xe^{x^{2}} + 4e^{4x}\] can we do something with the powers?

zepdrix (zepdrix):

yayyy good job \c:/ with the powers? like exponent rules? Hmm no, not really.

OpenStudy (anonymous):

Thank you zep and approach! :)

OpenStudy (xapproachesinfinity):

You're welcome